OpenStudy (anonymous):

Solve the exponental equation. (1/16)=64^4x-3 work if possible

Parth (parthkohli):

Observe:\[\dfrac{1}{16} = 16^{-1} = 4^{-2}\]and\[64^{4x - 3}=(4^3)^{4x - 3} =4^{12x - 9}\]

OpenStudy (anonymous):

im still confused, im sorry im horrible at this

Parth (parthkohli):

Notice that:\[4^{-2} = 4^{12x - 9}\]

OpenStudy (anonymous):

okay so what do i do after that?

Parth (parthkohli):

Then that means,\[-2 = 12x - 9\]

OpenStudy (anonymous):

okayy